let x be equal to a number another number is five less than twice x when the two numbers are multiplied the answer is 7 find two possible values for the number

Answer by checkley79(3341)

LET X BE THE FIRST NUMBER
LET 2X-5 BE THE SECOND NUMBER.
X(2X-5)=7
2X^2-5X-7=0
(2X-7)(X+1)=0
2X-7=0
2X=7
X=7/2
X=3.5 ANS.
2*X-5=2*3.5-5=7-5=2 ANS.
PROOF:
3.5*2=7
7=7